# Burger[Uncaptioned image]: Robust Graph Denoising-augmentation Fusion and Multi-semantic Modeling in Social Recommendation

### v2

@@ Abstract.
*Corresponding author.
copyright: nonejournalyear: 2025copyright: acmlicenseddoi: 10.1145/3652583.3658021isbn: 979-8-4007-0619-6/24/06
1. Introduction
With the development of social media, social recommendation as a hybrid recommendation system has attracted extensive attention. Previous research on human social behavior introduced two theories, i.e., the homophily theory (McPherson et al., 2001) and the social influence theory (Marsden and Friedkin, 1993). Based on the two theories, the core idea of social recommendation is to mine shared preference patterns among socially-connected users (Du et al., 2022). Early social recommendations (Yang et al., 2017; Jamali and Ester, 2010; Guo et al., 2015; Salakhutdinov and Mnih, 2007; Fang et al., 2014) are based on social relations, using trust propagation algorithms to predict preferences based on trust degrees, or leveraging friends’ interests for collaborative filtering. Later methods (Yuan et al., 2023; Liu et al., 2024; Tao et al., 2022; Mu et al., 2019) based on Graph Neural Networks (GNNs) construct social networks and user-item networks into graphs, and use graph convolution and graph attention mechanisms for message passing and feature aggregation to capture complex information, accurately mine interests, and improve recommendation quality.
Although GNNs-based social recommendations have shown their effectiveness, these methods overlook the fact that interest-irrelevant relations (Dunbar, 2010) in social networks can hinder performance of the model. On the one hand, users may establish connections with those who don’t have the same interests out of social politeness or on a whim. On the other hand, fake accounts and junk information are flooding. These accounts randomly add friends or follow others, which seriously disrupts the normal order of social networks. To alleviate the influence of interest-irrelevant relations, some studies (Yang et al., 2023, 2024) attempt to filter out these relations identified based on users’ item preferences. For example, GDMSR (Quan et al., 2023) leverages Transformer (Vaswani et al., 2017) to capture users’ item preferences, calculates the relation confidence between socially-connected users based on these preferences, and then iteratively deletes the edges of low relation confidence from original social graphs. DSL (Wang et al., 2023) adjusts the self-supervised signals by leveraging the similarity of users’ item preferences and filters out misleading social influences.
Figure 1. The mutual influence between the two semantic information.
@@ 2. Related Work
2.1. Social Recommendation based on GNNs.
Early social recommendation methods (Jamali and Ester, 2010; Ma et al., 2011; Guo et al., 2015; Yang et al., 2017) rely on matrix factorization, which fails to deeply capture user social-interest and item-interest connections. Graph neural networks address this by capturing higher-order, robust connections between users and items. For instance, DiffNet (Wu et al., 2019) and DiffNet++ (Wu et al., 2022) model social network topology, user interests, and social influence. Graphrec (Mu et al., 2019) uses graph attention networks to integrate user-item and social graph information. FBNE (Chen et al., 2022) captures high-order implicit relations, but these models are often biased by social noise.
2.2. Noise-resistant Social Recommendation
Noise-resistant social recommendation typically involves edge manipulation to clean noisy relations. These methods evaluate the connections between users by mining the preferences on the user-item bipartite graphs. For instance, existing works (Rey et al., 2022; Yoon et al., 2021) use probabilistic neighbor sampling to learn reliable neighboring nodes adaptively. GDMSR (Quan et al., 2023) employs sparsification for a cleaner social graph, while ESRF (Yu et al., 2020) utilizes GANs to generate denser social connections. On the other hand, some self-supervised models can also resist the interference of noise. They can extract more robust features by mining the self-supervised signals in the data, such as utilizing the contextual information of social networks, time series information, etc., thus effectively reducing the impact of noise on the recommendation results. For instance, Re4 (Zhang et al., 2022) and CML (Wei et al., 2022) leverage mutual interests to address noise, while DSL (Wang et al., 2023) aligns social and user-item network information through user sampling. Zhou et al. (Zhou et al., 2020) suggest enhancing sequential recommendation via self-supervised pre-training. MADM(Ma et al., 2024) alleviates social noise by maximizing the mutual information between users’ social representations and users’ item preference representations.
@@ 2.2. Noise-resistant Social Recommendation
3. Methodology
Figure 2. The overall framework of Burger model and the process of the (x+1)-th iteration.
@@ 3.1. Notations and Preliminaries
Task Description. According to the above definition, the tasks of the Burger model are as follows: given user-item interaction graph $G^{R}_{U,I}$ and user social graph $G^{S}_{U}$, the goal of Burger is to recommend unobserved items to target users based on the relevance of their embeddings extracted from these graphs.
3.2. Sliding Window based Social Tensor Construction
@@
(1) $\mathcal{T}=concatenate(G^{S_{-\tau+1}}_{U},G^{S_{-\tau+2}}_{U},\cdots,G^{S_{0}}_{U}).$
After the first iteration, Burger generates an enhanced slice $G^{S_{1}}_{U}$ based on the latest slice $G^{S_{0}}_{U}$ in the social tensor $\mathcal{T}$. The slice $G^{S_{1}}_{U}$ contains less noise and richer information. Inspired by (Liu et al., 2019), we propose to use sliding window mechanism (Liu et al., 2019) to update $\mathcal{T}$. The slice $G^{S_{1}}_{U}$ will be added to $\mathcal{T}$, while the oldest slice $G^{S_{-\tau+1}}_{U}$ in $\mathcal{T}$ will be removed. And the social tensor $\mathcal{T}$ is updated as follows:
(2) $\mathcal{T}=concatenate(G^{S_{-\tau+2}}_{U},G^{S_{-\tau+3}}_{U},\cdots,G^{S_{1}}_{U}).$
@@
(3) $\mathcal{T}=concatenate(G^{S_{x-\tau+2}}_{U},G^{S_{x-\tau+3}}_{U},\cdots,G^{S_{x+1}}_{U}).$
3.3. User Dual-preference Joint Mining
@@
3.3.1. users’ item preference mining.
Existing users’ item preference modeling largely relies on collaborative filtering methods. In these methods, graph convolutional networks (GCN) excel at capturing collaborative signals on the graph $G^{R}_{U,I}$, leading to improved recommendation performance. Taking the classic GCN model LightGCN(He et al., 2020) as an example, $\mathbf{E}_{U}\in\mathbb{R}^{m\times d}$ and $\mathbf{E}_{I}\in\mathbb{R}^{n\times d}$ represent the initial embedding matrices of users and items, respectively. The computation process of the $k$-th layer graph convolution in LightGCN is as follows:
(4) $\mathbf{e}_{u}^{(k+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{|\mathcal{N}_{u}|}\sqrt{|\mathcal{N}_{i}|}}\mathbf{e}_{i}^{(k)},$
(5) $\mathbf{e}_{i}^{(k+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{|\mathcal{N}_{i}|}\sqrt{|\mathcal{N}_{u}|}}\mathbf{e}_{u}^{(k)},$
@@
(6) $\mathbf{e}_{u}^{c}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{u}^{k},\ \ \mathbf{e}_{i}^{c}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{i}^{k},$
@@
(7) $\phi_{uv}=\mathbf{e}_{u}^{c}(\mathbf{e}_{v}^{c})^{\top}.$
3.3.2. users’ social preference mining.
@@
(8) $\mathcal{E}_{U}^{k+1}=\mathcal{T}\ast\mathcal{E}_{U}^{k}.$
@@
(9) $\mathcal{E}_{U}^{s}=\frac{1}{K+1}\sum_{k=0}^{K}\mathcal{E}_{U}^{k}.$
@@
(10) $\mathbf{E}_{U}^{s}=AGG(\mathbf{E}_{U}^{s,(1)},\cdots,\mathbf{E}_{U}^{s,(\tau)}),\ \ \ \mathbf{E}_{U}^{s,(\cdot)}\in\mathcal{E}_{U}^{s}.$
@@
(11) $\varphi_{uv}=\mathbf{e}_{u}^{s}(\mathbf{e}_{v}^{s})^{\top},\ \ \mathbf{e}_{\cdot}^{s}\in\mathbf{E}_{U}^{s},$
@@
3.4. Multi-semantic Modeling
After obtaining the interest similarity and social similarity of users, an intuitive idea is to leverage the recommendation loss to achieve better performance and model the mutual influence of users’ semantic behaviors. In most existing contrastive learning approaches (Hjelm et al., 2019), they maximize the agreement between positive samples while pushing negative pairs away in the embedding space, as shown below with derived gradients.
(12) $\frac{\partial L_{Infomax}}{\partial{\mathbf{e}}_{u}}=-{\mathbf{e}}_{v^{+}}+\sum_{v^{-}}{\mathbf{e}}_{v^{-}}\frac{\exp{\mathbf{e}}_{u}^{\top}{\mathbf{e}}_{v^{-}}}{\sum_{v^{-}}\exp{\mathbf{e}}_{u}^{\top}{\mathbf{e}}_{v^{-}}},$
@@ 3.4. Multi-semantic Modeling
(13) $\frac{\partial\mathcal{L}_{\Omega_{1}}}{\partial\mathbf{e}_{u}^{s}}=-\sigma(\phi_{uv^{+}})\mathbf{e}_{v^{+}}^{s}+\sigma(\phi_{uv^{-}})\mathbf{e}_{v^{-}}^{s}.$
@@
(14) $\frac{\partial\mathcal{L}_{\Omega_{2}}}{\partial\mathbf{e}_{u}^{c}}=-\sigma(\varphi_{uv^{+}})\mathbf{e}_{v^{+}}^{c}+\sigma(\varphi_{uv^{-}})\mathbf{e}_{v^{-}}^{c}.$
@@
(15) $=-\sigma(\phi_{uv^{+}})\varphi_{uv^{+}}+\sigma(\phi_{uv^{-}})\varphi_{uv^{-}}+C_{1},$
[…]
(16) $=-\sigma(\varphi_{uv^{+}})\phi_{uv^{+}}+\sigma(\varphi_{uv^{-}})\phi_{uv^{-}}+C_{2},$
(17) $\mathcal{L}_{\Omega_{1}}=\sum_{(u,v^{+},v^{-})}\max\left\{0,C_{1}-\sigma(\phi_{uv^{+}})\varphi_{uv^{+}}+\sigma(\phi_{uv^{-}})\varphi_{uv^{-}}\right\},$
(18) $\mathcal{L}_{\Omega_{2}}=\sum_{(u,v^{+},v^{-})}\max\left\{0,C_{2}-\sigma(\varphi_{uv^{+}})\phi_{uv^{+}}+\sigma(\varphi_{uv^{-}})\phi_{uv^{-}}\right\}.$
On the other hand, to achieve better recommendation performance, a classical BPR loss (Rendle et al., 2009) is utilized. Naturally, the likelihood of user $u$ interacting with item $i$ can be predicted as follows:
(19) $\rho_{ui}=\mathbf{e}_{u}^{c}(\mathbf{e}_{i}^{c})^{\top}.$
(20) $\mathcal{L}_{rec}=-\sum_{(u,i^{+},i^{-})}ln\ \sigma(\rho_{ui^{+}}-\rho_{ui^{-}}).$
(21) $\mathcal{L}_{soc}=-\sum_{(u,v^{+},v^{-})}ln\ \sigma(\varphi_{uv^{+}}-\varphi_{uv^{-}}).$
(22) $\mathcal{L}=\mathcal{L}_{rec}+\lambda_{1}\mathcal{L}_{soc}+\alpha\mathcal{L}_{\Omega_{1}}+\beta\mathcal{L}_{\Omega_{2}}+\lambda_{2}(||\mathbf{E}_{U}||+||\mathbf{E}_{I}||),$
@@
3.5. Graph Denoising-augmentation Fusion
Merely modeling the mutual influence of semantic information is unreliable, especially $\mathcal{L}_{\Omega_{2}}$, because social noise can cause the user pairs corresponding to the noise to have their distances forcibly coordinated in the semantic space $\Omega_{2}$, which can lead to suboptimal performance. Therefore, it is necessary to remove social noise. In addition, some negative users who are not yet observed by an anchor user, and there may exist potential social connections, such as some users who may make friends with the anchor user in the near future. And these negative users are classified as potential friends or potential non-friends for an anchor user. Therefore, it is necessary to mine potential friends with the same item preferences, which can improve the generalization ability and performance of the model. Based on the two ideas, this module can be divided into two steps: First, identify potential friends from those not observed by the anchor user; second, identify potential friends who have shared item preferences with the anchor user and replace noisy users. Specifically, for user $u$, the set of $u$’s friends observed in the social graph to be enhanced is denoted as $\mathcal{B}_{u}$, i.e., the set of positive users and $v^{+}\in\mathcal{B}_{u}$. The set of friends not observed by the user $u$ can then be easily obtained: $\mathcal{C}_{u}=U-\mathcal{B}_{u}-u$, i.e., the set of negative users. An intuitive idea is to identify potential friends from $\mathcal{C}_{u}$. However, identifying potential friends is non-trivial due to a lack of ground-truth labels.
For this problem, we attribute it to negative sample sampling. Specifically, in the unobserved set, one type of negative sample is one that anchor users will not interact with, while another type of negative sample may be one that anchor users will interact with; The latter is a pseudo negative sample, while the former is a true negative sample. Inspired by existing work (Liu and Wang, 2023), we attempt to use Bayesian posterior probability to calculate the probability of unobserved users as pseudo negative samples, and the set of users with the higher probability is selected as the identified potential friends. Specifically, for user $v^{-}\in\mathcal{C}_{u}$, the probability $pf_{u}(v^{-})$ of being a potential friend of $u$ can be computed as follows:
(23) $pf_{u}(v^{-})=\frac{F(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})}{1-F(\varphi_{uv^{-}})-P_{\mathfrak{f}}(v^{-})+2F(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})},$
where $F(\varphi_{uv^{-}})=\frac{|\{w|\forall w\in\mathcal{C}_{u},\varphi_{uw}<\varphi_{uv^{-}}\}|}{\left|\mathcal{C}_{u}\right|}$, and $P_{\mathfrak{f}}(v^{-})$ represents the prior probability that $v^{-}$ is a potential friend. The prior probability can be determined using exposure rates and other information. The derivation process of this Eq.23 is discussed in section 4.3.1.

### v3

# Burger[Uncaptioned image]: Robust Graph Denoising-augmentation Fusion and Multi-semantic Modeling for Social Recommendation

@@ Abstract
I Introduction
With the development of social media, social recommendation as a hybrid recommendation system has attracted extensive attention. Previous research on human social behavior introduced two theories, i.e., the homophily theory [1] and the social influence theory [2]. The former refers to the fact that people with common interests are more likely to establish social relations, while the latter means that social relations can influence individuals’ interests and behaviors.
Based on the two theories, the core idea of social recommendation is to mine shared preference patterns among socially-connected users [3]. Early social recommendations [4, 5, 6, 7, 8, 9, 10] are based on social relations, using trust propagation algorithms to predict preferences based on trust degrees, or leveraging friends’ interests for collaborative filtering. Later methods [11, 12, 13, 14] based on Graph Neural Networks (GNNs) construct social networks and user-item networks into graphs, and use graph convolution and graph attention mechanisms for message passing and feature aggregation to capture complex information, accurately mine interests, and improve recommendation quality.
Figure 1: A toy example illustrating the mutual influence between the two semantic information.
Although GNNs-based social recommendations have shown their effectiveness, these methods overlook the fact that interest-irrelevant relations [15] in social networks can hinder performance of the model. On the one hand, users may establish connections with those who don’t have the same interests out of social politeness or on a whim. On the other hand, fake accounts and junk information are flooding. These accounts randomly add friends or follow others, which seriously disrupts the normal order of social networks. To alleviate the influence of interest-irrelevant relations, some studies [16, 17] attempt to filter out these relations identified based on users’ item preferences. For example, GDMSR [18] leverages Transformer [19] to capture users’ item preferences, calculates the relation confidence between socially-connected users based on these preferences, and then iteratively deletes the edges of low relation confidence from original social graphs. DSL [20] adjusts the self-supervised signals by leveraging the similarity of users’ item preferences and filters out misleading social influences.
@@ II Related Work
II-A Social Recommendation based on GNNs
Early social recommendation methods [5, 21, 6, 4] rely on matrix factorization, which fails to deeply capture user social-interest and item-interest connections. Graph neural networks address this by capturing higher-order, robust connections between users and items. For instance, DiffNet [22] and DiffNet++ [23] model social network topology, user interests, and social influence. Graphrec [14] uses graph attention networks to integrate user-item and social graph information. FBNE [24] captures high-order implicit relations, but these models are often biased by social noise.
II-B Noise-resistant Social Recommendation
Noise-resistant social recommendation typically involves edge manipulation to clean noisy relations. These methods evaluate the connections between users by mining the preferences on the user-item bipartite graphs. For instance, existing works [25, 26] use probabilistic neighbor sampling to learn reliable neighboring nodes adaptively. GDMSR [18] employs sparsification for a cleaner social graph, while ESRF [27] utilizes GANs to generate denser social connections. On the other hand, some self-supervised models can also resist the interference of noise. They can extract more robust features by mining the self-supervised signals in the data, such as utilizing the contextual information of social networks, time series information, etc., thus effectively reducing the impact of noise on the recommendation results. For instance, Re4 [28] and CML [29] leverage mutual interests to address noise, while DSL [20] aligns social and user-item network information through user sampling. Zhou et al. [30] suggest enhancing sequential recommendation via self-supervised pre-training. MADM[31] alleviates social noise by maximizing the mutual information between users’ social representations and users’ item preference representations.
@@ II-B Noise-resistant Social Recommendation
III Methodology
Figure 2: The overall framework of Burger model and the process of the (x+1)-th iteration. The Burger model consists of four modules. Module 1 represents sliding window based social tensor construction module, Module 2 represents user dual-preference joint mining module, Module 3 represents multi-semantic modeling module and Module 4 represents graph denoising-augmentation fusion module.
@@ III-A Notations and Preliminaries
According to the above definition, the tasks of the Burger model are as follows: given user-item interaction graph $G^{R}_{U,I}$ and user social graph $G^{S}_{U}$, the goal of Burger is to recommend unobserved items to target users based on the relevance of their embeddings extracted from these graphs.
III-B Sliding Window based Social Tensor Construction
@@
$\mathcal{T}=concatenate(G^{S_{-\tau+1}}_{U},G^{S_{-\tau+2}}_{U},\cdots,G^{S_{0}}_{U}).$ (1)
After the first iteration, Burger generates an enhanced slice $G^{S_{1}}_{U}$ based on the latest slice $G^{S_{0}}_{U}$ in the social tensor $\mathcal{T}$. The slice $G^{S_{1}}_{U}$ contains less noise and richer information. Inspired by [32], we propose to use sliding window mechanism [32] to update $\mathcal{T}$. The slice $G^{S_{1}}_{U}$ will be added to $\mathcal{T}$, while the oldest slice $G^{S_{-\tau+1}}_{U}$ in $\mathcal{T}$ will be removed. And the social tensor $\mathcal{T}$ is updated as follows:
$\mathcal{T}=concatenate(G^{S_{-\tau+2}}_{U},G^{S_{-\tau+3}}_{U},\cdots,G^{S_{1}}_{U}).$ (2)
@@
$\mathcal{T}=concatenate(G^{S_{x-\tau+2}}_{U},G^{S_{x-\tau+3}}_{U},\cdots,G^{S_{x+1}}_{U}).$ (3)
III-C User Dual-preference Joint Mining
@@
III-C1 users’ item preference mining
Existing users’ item preference modeling largely relies on collaborative filtering methods. In these methods, graph convolutional networks (GCN) excel at capturing collaborative signals on the graph $G^{R}_{U,I}$, leading to improved recommendation performance. Taking the classic GCN model LightGCN[33] as an example, $\mathbf{E}_{U}\in\mathbb{R}^{m\times d}$ and $\mathbf{E}_{I}\in\mathbb{R}^{n\times d}$ represent the initial embedding matrices of users and items, respectively. The computation process of the $k$-th layer graph convolution in LightGCN is as follows:
$\mathbf{e}_{u}^{(k+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{|\mathcal{N}_{u}|}\sqrt{|\mathcal{N}_{i}|}}\mathbf{e}_{i}^{(k)},$ (4)
$\mathbf{e}_{i}^{(k+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{|\mathcal{N}_{i}|}\sqrt{|\mathcal{N}_{u}|}}\mathbf{e}_{u}^{(k)},$ (5)
@@
$\mathbf{e}_{u}^{c}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{u}^{k},\ \ \mathbf{e}_{i}^{c}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{i}^{k},$ (6)
@@
$\phi_{uv}=\mathbf{e}_{u}^{c}(\mathbf{e}_{v}^{c})^{\top}.$ (7)
III-C2 users’ social preference mining
@@
$\mathcal{E}_{U}^{k+1}=\mathcal{T}\ast\mathcal{E}_{U}^{k}.$ (8)
@@
$\mathcal{E}_{U}^{s}=\frac{1}{K+1}\sum_{k=0}^{K}\mathcal{E}_{U}^{k}.$ (9)
@@
$\mathbf{E}_{U}^{s}=AGG(\mathbf{E}_{U}^{s,(1)},\cdots,\mathbf{E}_{U}^{s,(\tau)}),\ \ \ \mathbf{E}_{U}^{s,(\cdot)}\in\mathcal{E}_{U}^{s}.$ (10)
@@
$\varphi_{uv}=\mathbf{e}_{u}^{s}(\mathbf{e}_{v}^{s})^{\top},\ \ \mathbf{e}_{\cdot}^{s}\in\mathbf{E}_{U}^{s},$ (11)
@@
III-D Multi-semantic Modeling
After obtaining the interest similarity and social similarity of users, an intuitive idea is to leverage the recommendation loss to achieve better performance and model the mutual influence of users’ semantic behaviors. In most existing contrastive learning approaches [34], they maximize the agreement between positive samples while pushing negative pairs away in the embedding space, as shown below with derived gradients.
$\frac{\partial L_{Infomax}}{\partial{\mathbf{e}}_{u}}=-{\mathbf{e}}_{v^{+}}+\sum_{v^{-}}{\mathbf{e}}_{v^{-}}\frac{\exp{\mathbf{e}}_{u}^{\top}{\mathbf{e}}_{v^{-}}}{\sum_{v^{-}}\exp{\mathbf{e}}_{u}^{\top}{\mathbf{e}}_{v^{-}}},$ (12)
@@ III-D Multi-semantic Modeling
$\frac{\partial\mathcal{L}_{\Omega_{1}}}{\partial\mathbf{e}_{u}^{s}}=-\sigma(\phi_{uv^{+}})\mathbf{e}_{v^{+}}^{s}+\sigma(\phi_{uv^{-}})\mathbf{e}_{v^{-}}^{s}.$ (13)
@@
$\frac{\partial\mathcal{L}_{\Omega_{2}}}{\partial\mathbf{e}_{u}^{c}}=-\sigma(\varphi_{uv^{+}})\mathbf{e}_{v^{+}}^{c}+\sigma(\varphi_{uv^{-}})\mathbf{e}_{v^{-}}^{c}.$ (14)
@@
$\int\frac{\partial\mathcal{L}_{\Omega_{1}}}{\partial\mathbf{e}_{u}^{s}}d\mathbf{e}_{u}^{s}$ $=\int\left[-\sigma(\phi_{uv^{+}})\mathbf{e}_{v^{+}}^{s}+\sigma(\phi_{uv^{-}})\mathbf{e}_{v^{-}}^{s}\right]d\mathbf{e}_{u}^{s}$
$=-\sigma(\phi_{uv^{+}})\varphi_{uv^{+}}+\sigma(\phi_{uv^{-}})\varphi_{uv^{-}}+C_{1},$ (15)
$\int\frac{\partial\mathcal{L}_{\Omega_{2}}}{\partial\mathbf{e}_{u}^{c}}d\mathbf{e}_{u}^{c}$ $=\int\left[-\sigma(\varphi_{uv^{+}})\mathbf{e}_{v^{+}}^{c}+\sigma(\varphi_{uv^{-}})\mathbf{e}_{v^{-}}^{c}\right]d\mathbf{e}_{u}^{c}$
$=-\sigma(\varphi_{uv^{+}})\phi_{uv^{+}}+\sigma(\varphi_{uv^{-}})\phi_{uv^{-}}+C_{2},$ (16)
$\mathcal{L}_{\Omega_{1}}=\sum_{(u,v^{+},v^{-})}\max\left\{0,C_{1}-\sigma(\phi_{uv^{+}})\varphi_{uv^{+}}+\sigma(\phi_{uv^{-}})\varphi_{uv^{-}}\right\}$ (17)
$\mathcal{L}_{\Omega_{2}}=\sum_{(u,v^{+},v^{-})}\max\left\{0,C_{2}-\sigma(\varphi_{uv^{+}})\phi_{uv^{+}}+\sigma(\varphi_{uv^{-}})\phi_{uv^{-}}\right\}$ (18)
On the other hand, to achieve better recommendation performance, a classical BPR loss [35] is utilized. Naturally, the likelihood of user $u$ interacting with item $i$ can be predicted as follows:
$\rho_{ui}=\mathbf{e}_{u}^{c}(\mathbf{e}_{i}^{c})^{\top}.$ (19)
$\mathcal{L}_{rec}=-\sum_{(u,i^{+},i^{-})}ln\ \sigma(\rho_{ui^{+}}-\rho_{ui^{-}}).$ (20)
$\mathcal{L}_{soc}=-\sum_{(u,v^{+},v^{-})}ln\ \sigma(\varphi_{uv^{+}}-\varphi_{uv^{-}}).$ (21)
@@
III-E Graph Denoising-augmentation Fusion
Merely modeling the mutual influence of semantic information is unreliable, especially $\mathcal{L}_{\Omega_{2}}$, because social noise can cause the user pairs corresponding to the noise to have their distances forcibly coordinated in the semantic space $\Omega_{2}$, which can lead to poor performance. Therefore, it is necessary to remove social noise. In addition, only focusing on discarding irrelevant social connections makes sparse social networks even sparser, meaning social representations may not be sufficient for recommendation. Therefore, it is necessary to mine potential friends with the same item preferences, which can improve the generalization ability and performance of the model. Based on the two ideas, this module can be divided into two steps: First, identify potential friends from those not observed by the anchor user; second, identify potential friends who have shared item preferences with the anchor user and replace noisy users. Specifically, for user $u$, the set of $u$’s friends observed in the social graph to be enhanced is denoted as $\mathcal{B}_{u}$, i.e., the set of positive users and $v^{+}\in\mathcal{B}_{u}$. The set of friends not observed by the user $u$ can then be easily obtained: $\mathcal{C}_{u}=U-\mathcal{B}_{u}-u$, i.e., the set of negative users. An intuitive idea is to identify potential friends from $\mathcal{C}_{u}$. However, identifying potential friends is non-trivial due to a lack of ground-truth labels.
For this problem, we attribute it to negative sample sampling. Specifically, in the unobserved set, one type of negative sample is one that anchor users will not interact with, while another type of negative sample may be one that anchor users will interact with; The latter is a pseudo negative sample, while the former is a true negative sample. Inspired by existing work [36], we attempt to use Bayesian posterior probability to calculate the probability of unobserved users as pseudo negative samples, and the set of users with the higher probability is selected as the identified potential friends. Specifically, for user $v^{-}\in\mathcal{C}_{u}$, the probability $pf_{u}(v^{-})$ of being a potential friend of $u$ can be computed as follows:
$pf_{u}(v^{-})=\frac{F(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})}{1-F(\varphi_{uv^{-}})-P_{\mathfrak{f}}(v^{-})+2F(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})},$ (23)
where $F(\varphi_{uv^{-}})=\frac{|\{w|\forall w\in\mathcal{C}_{u},\varphi_{uw}<\varphi_{uv^{-}}\}|}{\left|\mathcal{C}_{u}\right|}$, and $P_{\mathfrak{f}}(v^{-})$ represents the prior probability that $v^{-}$ is a potential friend. The prior probability can be determined using exposure rates and other information. The derivation process of this Eq.23 is discussed in section IV-C1.